OpenStudy (anonymous):

Find the length of side a in the right triangle, if b=12 and c=15

OpenStudy (anonymous):

9

OpenStudy (anonymous):

a-squared + b-squared = c-squared. a-squared + 144 = 225 sqrt81 = 9

OpenStudy (anonymous):

the anser is 9 bro

OpenStudy (anonymous):

\[\left( 15 \right)^{2} = a ^{2} + \left( 12 \right)^{2}\] \[225 = a ^{2} + 144\] \[\sqrt[2]81 = \sqrt[2] a ^{2}\] a = 9